Determine the Z parameters of the network shown below and check the symmetry and reciprocity of the network. Also find the h parameters of the network. 


1
Expert's answer
2021-07-12T13:27:23-0400

We know that

"V_1=z_{11}I_1+z_{12}I_2"

"V_2=z_{21}I_1+z_{22}I_2"

"h_{11}=\frac{V_1}{I_1},V_2=0"

"h_{21}=\frac{I_2}{I_1},V_2=0"

"V_2=0"

"z_{21}I_1+z_{22}I_2=0"

"h_{21}=\frac{I_2}{I_1}=-\frac{z_{21}}{z_{22}}"

"h_{11}=\frac{z_{11}z_{22}-z_{12}z_{21}}{z_{22}}=\frac{∆_z}{z_{11}}"

"V_1=z_{12}I_2"


"V_2=z_{22}I_2"

"h_{12}=\frac{V_1}{V_2},I_1=0"

"h_{22}=\frac{I_2}{V_2},I_1=0"

"h_{12}=\frac{V_1}{V_2}=\frac{z_{12}}{z_{22}}"

"h_{22}=\frac{I_2}{V_2}=\frac{1}{z_{22}}"

"\begin{bmatrix}
 h_{11} & h_{12} \\
 h_{13} & h_{14}
\end{bmatrix}" "=\begin{bmatrix}
 \frac{∆_z}{z_{22}} & \frac{z_{12}}{z_22}\\
 -\frac{z_{21}}{z_{22}} & \frac{1}{z_{22}}
\end{bmatrix}"